OpenStudy (anonymous):

Find dy/dx. y= cos(t^2)dt on the integral from sqrt(x) to pi/3

OpenStudy (anonymous):

u have to dy/dx of \[\int\limits_{\sqrt{x}}^{\pi/3} y dx\]

OpenStudy (anonymous):

where y= cos(t^2) dt

OpenStudy (anonymous):

\[\int\limits_{g(x)}^{h(x)} f(x) dx =y\] then dy/dx= f(g(x)) g'(x) - f(h(x)) h'(x)

OpenStudy (anonymous):

Thank you, I'm getting confused with what to do with the cos(pi/3)^2, the cos of pi/3 is 1/2, so am I sqauring the 1/2?

OpenStudy (anonymous):

sorry ...dy/dx= f(h(x)) h'(x) - f(g(x)) g'(x)

OpenStudy (anonymous):

here h(x)= pi/3 so h'(x)=0

OpenStudy (anonymous):

f(h(x))= cos[ (pi/3)^2) ]

OpenStudy (anonymous):

so use calculator for finding cos[ (pi/3)^2) ]

OpenStudy (anonymous):

but h'(x) =0 so f(h(x)) h'(x)=0

OpenStudy (anonymous):

does it make sense?

OpenStudy (anonymous):

So is my answer (cosx)/(2sqrtx)? or 1/2* cosx/sqrtx?

OpenStudy (anonymous):

- cos(x) [(1/ {2sqrt(x)} ]

OpenStudy (anonymous):

Oh thanks, I forgot to switch the h(x) and g(x)

OpenStudy (anonymous):

\[\int\limits\limits_{g(x)}^{h(x)} f(x) dx =y\] then dy/dx= f(h(x)) h'(x) - f(g(x)) g'(x)

OpenStudy (anonymous):

Thank you for your help!

OpenStudy (anonymous):

:)